OpenStudy (anonymous):

A sample of gas with a mass of 8.40 g occupies a volume of 4.50 L at 298 K and 1.10 atm. What is the molar mass of the gas? Show all work used to find your answer. *Best answer gets medal, fan and testimony*

OpenStudy (aaronq):

Use the ideal gas law, substitute in the definition of moles, solve for M, molar mass. Ideal gas law: \(\sf PV=nRT\) moles: \(\sf n=\dfrac{m}{M}\) \(\sf \huge PV=\dfrac{m}{M}RT\rightarrow M=\dfrac{mRT}{PV}\)

OpenStudy (anonymous):

May you help me plug in the variables?

OpenStudy (aaronq):

I can check it for you if you give it a try

OpenStudy (anonymous):

Ok. Give me a second.

OpenStudy (anonymous):

I have: 1.10*4.50 L = n(0.08206atm)(298K) And that's all I can do. Unfortunately.

OpenStudy (aaronq):

that's right! you can now solve for n, then divide the mass 8.4 g by the moles.

OpenStudy (aaronq):

the mass is given in the question, it's 8.4 grams

OpenStudy (anonymous):

So it would be: 1.10 atm * 298 K = 24.45388 8.4 g / 24.45388 = 0.3435

OpenStudy (aaronq):

not quite, first solve this equation: 1.10*4.50 L = n(0.08206atm)(298K) divide the mass by the answer (moles)

OpenStudy (anonymous):

4.95 = n(0.08206*298) 4.95 = n(24.45388) 4.95/24.45388 = 0.202421... n = 0.202421

OpenStudy (aaronq):

okay, looks good! now divide the mass by n

OpenStudy (anonymous):

244.5398

OpenStudy (aaronq):

There you go

OpenStudy (anonymous):

WOOHOO! How would I properly write with sig figs?

OpenStudy (aaronq):

all the numbers used have 3 sig figs, so make the answer have 3 figures.

OpenStudy (anonymous):

244. Thank you so much! You're a life saver :D

OpenStudy (aaronq):

that's not quite right. look at the fourth digit, you have to round up

OpenStudy (anonymous):

Oh... 245, correct?

OpenStudy (aaronq):

yeah. you should give this a look https://www.physics.uoguelph.ca/tutorials/sig_fig/SIG_dig.htm

OpenStudy (anonymous):

Will do.

OpenStudy (anonymous):

Got it. Thanks :)

OpenStudy (aaronq):

no problem!

OpenStudy (anonymous):

@aaronq there is problem in this question

OpenStudy (aaronq):

whats the problem?

OpenStudy (anonymous):

should the answer be around 41 to 42 because we are dividing 8.4 by n

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

you're right, the answer is 41.4974933333333333 I guess this person plugged the numbers into the calc wrong

OpenStudy (anonymous):

oh thank so much

OpenStudy (aaronq):

np!

OpenStudy (aaronq):

thanks for double checking lol

OpenStudy (anonymous):

np!